OpenStudy (anonymous):

what are the prime factors of 144

OpenStudy (anonymous):

11 * 2 * 2

OpenStudy (anonymous):

wrong answer^^

OpenStudy (anonymous):

(3*2^2)^2

OpenStudy (anonymous):

damn >.>

OpenStudy (anonymous):

well, i tried.

OpenStudy (anonymous):

that is 3*3*2*2*2*2

OpenStudy (anonymous):

thank ya.

OpenStudy (anonymous):

To find the various prime factors of 144. just express break down 144. Like in the following, 144= 2 * 72 = 2* 2 * 36 = 2* 2 * 2* 18 = 2 * 2* 2* 2* 9 = 2* 2* 2* 2* 3*3 = 2^4 * 3^2 Just keep dividing to break it down step by step.

OpenStudy (anonymous):

dg123 thats the answer:) thanks!

OpenStudy (anonymous):

anytime bro :)